### v1

# DAASH: A Meta-Attack Framework for Synthesizing Effective and Stealthy Adversarial Examples

Abstract
Numerous techniques have been proposed for generating adversarial examples in white-box settings under strict $\ell_{p}$-norm constraints. However, such norm-bounded examples often fail to align well with human perception, and only recently have a few methods begun specifically exploring perceptually aligned adversarial examples. Moreover, it remains unclear whether insights from $\ell_{p}$-constrained attacks can be effectively leveraged to improve perceptual efficacy. In this paper, we introduce DAASH, a fully differentiable meta-attack framework that generates effective and perceptually aligned adversarial examples by strategically composing existing $\ell_{p}$-based attack methods. DAASH operates in a multi-stage fashion: at each stage, it aggregates candidate adversarial examples from multiple base attacks using learned, adaptive weights and propagates the result to the next stage. A novel meta-loss function guides this process by jointly minimizing misclassification loss and perceptual distortion, enabling the framework to dynamically modulate the contribution of each base attack throughout the stages. We evaluate DAASH on adversarially trained models across CIFAR-10, CIFAR-100, and ImageNet. Despite relying solely on $\ell_{p}$-constrained based methods, DAASH significantly outperforms state-of-the-art perceptual attacks such as AdvAD—achieving higher attack success rates (e.g., 20.63% improvement) and superior visual quality, as measured by SSIM, LPIPS, and FID (improvements of $\approx 11$, 0.015, and 5.7, respectively). Furthermore, DAASH generalizes well to unseen defenses, making it a practical and strong baseline for evaluating robustness without requiring handcrafted adaptive attacks for each new defense.
1 Introduction
Adversarial examples against deep neural networks (DNNs), which add imperceptible perturbations to clean images to induce misclassification (targeted/untargeted) Goodfellow et al. (2015a), have attracted significant attention from the research community. Various attack methods, including FGSM Goodfellow et al. (2015b), PGD Madry et al. (2018b), and CW Carlini and Wagner (2017), as well as adaptive evaluation frameworks such as AutoAttack Croce et al. (2020), have been developed to benchmark both defended and undefended models. However, the majority of these approaches prioritize imperceptibility by constraining perturbations within a small $\ell_{p}$-norm ball, with $\ell_{2}$ and $\ell_{\infty}$ norms being the most common ones.
Yet, minimizing $\ell_{p}$-norm does not always yield high perceptual similarity to the human visual system Zhang et al. (2018); Laidlaw and Feizi (2020); Sharif et al. (2018), despite imperceptibility being a fundamental design goal of stealthy adversarial examples. This disconnect has motivated the adoption of perceptually-aligned similarity metrics such as Structural Similarity Index (SSIM) Wang et al. (2004), Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018), and Fréchet Inception Distance (FID) Heusel et al. (2017) to better assess visual quality Li et al. (2024); Chen et al. (2024b). However, attacks optimized under these perceptual metrics often suffer from reduced success rates—particularly against robust, defended models (see Table 1)—highlighting a critical trade-off and leaving significant room for improvement.
Designing effective attacks against a wide range of models, especially hardened ones, remains non-trivial and often requires deep, manual effort—an observation consistently echoed in the literature on $\ell_{p}$-norm attacks Carlini and Wagner (2017); Athalye et al. (2018); Croce and Hein (2020b). This complexity intensifies when attempting to simultaneously preserve high perceptual fidelity, thus underscoring the need for systematic and automated strategies that can generalize across models while aligning better with human vision.
Figure 1: Comparative visualization of adversarial perturbations across multiple attack methods (including DAASH).
Given the extensive body of work on $\ell_{p}$-bounded attacks, including those targeting robust models, it is natural to ask:
Can we leverage the strengths of these existing attacks to compose stronger adversarial examples that are also more perceptually aligned?
Our key insight is that although individual $\ell_{p}$-norms (e.g., $\ell_{2}$, $\ell_{\infty}$) do not fully capture human perception, they each correlate with different perceptual characteristics (e.g., contrast, texture, edge disruption) Sharif et al. (2018); Laidlaw and Feizi (2020). Therefore, by combining attack strategies optimized under different norms in a principled manner, we may achieve not only improved attack success but also better perceptual quality, particularly when targeting defended models where individual attack methods may fall short. Specifically, we propose DAASH (Differential Adversarial Attack SearcH)—a fully differentiable, multi-stage framework that learns to compose multiple $\ell_{p}$-norm–constrained base attacks using soft attention. DAASH operates in two stages: (1) it computes a soft attention distribution over adversarial examples generated from individual candidate attacks (e.g., PGD, CW) to weight and compose the final adversarial example; (2) it iteratively refines the perturbations across multiple stages by feeding the output from one stage as the input to the next. This iterative chaining is designed to escape poor local minima inherent to the non-convex landscape of adversarial optimization.
The attention weights are optimized end-to-end through a meta-loss that jointly considers (i) attack success (via minimizing the confidence score of the ground-truth class) and (ii) perceptual similarity (via maximizing SSIM). This allows the framework to dynamically adjust the influence of each candidate attack based on the target model. Unlike existing methods that are rigid in balance between imperceptibility (from human perception) and attack success, DAASH learns to trade off these objectives adaptively across different models and datasets.
Empirically, DAASH achieves a significant improvement in both attack success rate (ASR) and perceptual fidelity compared to the current state-of-the-art. On robust models trained on CIFAR-100 such as Cui et al. (2024), DAASH reaches an average ASR of 99.97% with an SSIM score of 94.43, significantly outperforming state-of-the-art perceptual attacks such as AdvAD Li et al. (2024) (ASR: 79.14%, SSIM: 83.18) and DiffAttack Chen et al. (2024b) (ASR: 77.24%, SSIM: 91.16). Similar improvements are observed in other experimental settings.
Fig. 1 presents qualitative comparisons, further illustrating the superiority of DAASH in generating perceptually plausible yet highly effective adversarial examples.
Our contributions are summarized as follows:
  • We propose DAASH, a fully differentiable, multi-stage framework that learns to combine $\ell_{p}$-norm–constrained base attacks using soft attention.
  • We introduce a novel meta-loss that decouples attack effectiveness and human perceptual quality from the constraints of individual norm-based attacks, enabling better identification of attack strengths across models.
  • We validate DAASH across multiple datasets (CIFAR-10, CIFAR-100, ImageNet), seven robust models, and four post-processing defenses demonstrating substantial improvements over prior perceptual attacks.
2 Background and Threat Model
Adversarial Examples. Given a deep neural network based classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$, which maps inputs to class probabilities, and a clean sample $(x,y)$ where $x\in\mathcal{X}$ and $y\in\mathcal{Y}$, an untargeted adversarial example $\boldsymbol{x}_{a}$ is defined as:
$\arg\max_{i}f(\boldsymbol{x}_{a})_{i}\neq y\quad\text{subject to}\quad d(\boldsymbol{x},\boldsymbol{x}_{a})\leq\epsilon$ (1)
Imperceptibility constraints. In the equation above, $d(\cdot,\cdot)$ represents a distance metric and $\epsilon$ defines the perturbation budget. $d(\cdot,\cdot)$ measures the imperceptibility of the perturbation, and the commonly used metrics include $\ell_{2}$ norm Carlini and Wagner (2017) as well as the $\ell_{\infty}$-norm Carlini and Wagner (2017). These $\ell_{p}$-norm constraints may not always align with human perception, and some alternative metrics such as the Structural Similarity Index (SSIM) Wang et al. (2004) is used. SSIM is shown to better capture the luminance, contrast, and structural similarities between two images $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ and is computed as below:
$\text{SSIM}(\boldsymbol{x}_{1},\boldsymbol{x}_{2})=\frac{(2\mu_{x_{1}}\mu_{x_{2}}+c_{1})(2\sigma_{x_{1}x_{2}}+c_{2})}{(\mu_{x_{1}}^{2}+\mu_{x_{2}}^{2}+c_{1})(\sigma_{x_{1}}^{2}+\sigma_{x_{2}}^{2}+c_{2})}$ (2)
The terms $\mu_{x_{1}}$ and $\mu_{x_{2}}$ represent the mean pixel intensities of $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, respectively, while $\sigma_{x_{1}}^{2}$ and $\sigma_{x_{2}}^{2}$ denote their variances. The term $\sigma_{x_{1}x_{2}}$ corresponds to the covariance between $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$. The constants $c_{1}$ and $c_{2}$ are used to stabilize the division when the denominator is close to zero.
Besides SSIM, other perceptual similarity metrics that better align with human perception include the Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018) and the Fréchet Inception Distance (FID) Heusel et al. (2017). In this paper, we use SSIM as the primary metric to optimize for imperceptibility, but we find that using LPIPS and FID yields similar results in practice.
Threat model. In this paper, we focus on the white-box setting, where the attacker is assumed to have complete access to $f(\cdot)$, including the model architecture and parameters. In this setting, untargeted adversarial examples, as defined in Eqn. (1), can be generated via backpropagation using well-defined loss functions with respect to the input $\boldsymbol{x}$. The white-box assumption is widely adopted in the extensive literature Carlini and Wagner (2017); Madry et al. (2018a); Croce and Hein (2020b).
3 Related Work
Classical approaches for generating adversarial examples include FGSM Goodfellow et al. (2015a), I-FGSM Kurakin et al. (2016), and PGD Madry et al. (2018a), all of which strictly enforce a perturbation budget $\epsilon$ measured in the $\ell_{\infty}$-norm. In contrast, the CW attack Carlini and Wagner (2017) softly enforces an $\ell_{2}$-norm perturbation constraint by incorporating the perturbation magnitude directly into the optimization objective. Closer to our approach are AutoAttack Croce and Hein (2020b) and Composite Attacks Mao et al. (2021). AutoAttack employs an ensemble of a few selected attack methods, while Composite Attacks combine multiple attacks using genetic algorithms. However, these approaches still operate under conventional $\ell_{p}$-norm constraints and rely on simplistic or heuristic strategies for combining base attacks. In contrast, our proposed method DAASH introduces an optimization-based meta-framework to adaptively compose base attacks for improved attack effectiveness and imperceptibility for human eyes.
Recognizing the limitations of norm-constrained attacks in modeling human perception, recent works have shifted toward leveraging perceptual similarity metrics to guide adversarial example generation. For instance, SSAH Luo et al. (2022) operates in the frequency domain to introduce perturbations in less perceptible high-frequency components, while PerC-AL Zhao et al. (2020) optimizes perceptual color distance, though it relies on manual hyperparameter tuning. More recent approaches exploit diffusion models by injecting perturbations into their embedding spaces, thereby influencing the denoising trajectory to produce adversarial outputs. Representative examples include DiffAttack Chen et al. (2024a), AdvAD Li et al. (2024), and DiffPGD Xue et al. (2023). Despite their innovations, these perceptual attacks are often designed as standalone strategies with limited adaptability, making them less effective across diverse target models. As demonstrated in Table 1, such methods frequently underperform in both attack success and imperceptibility. By contrast, DAASH builds on the extensive literature of norm-constrained attacks—known for their compatibility with a wide range of target models—and combines them in a unified framework that enhances the effectiveness and stealth of the resulting adversarial examples.
4 DAASH Attack Design
Next, we describe the high-level motivation behind the design of DAASH and then introduce the details more formally.
Figure 2: An overview of the proposed Differentiable Adversarial Attack Search (DAASH) methodology.
4.1 Design motivation
Our attack design is motivated by two key observations: (1) Norm-constrained attacks sometimes exhibit complementary strengths and weaknesses Croce and Hein (2020b). So, combining them can yield better performance, especially in under-explored settings of generating adversarial examples beyond $\ell_{p}$-norm constraints; (2) Adversarial example generation is a non-convex optimization process and is prone to getting stuck in poor local optima Madry et al. (2018a).
Regarding Observation 1, the most natural way to combine different attacks is to take a weighted average of the adversarial examples generated by each individual attack. While a linear combination is simple and practical, it introduces two key challenges: (a) How to determine optimal weights for each base attack, and (b) How to ensure that the resulting adversarial example better aligns with human perception rather than just staying within the $\ell_{p}$ constraints.
To address these, we make the combination weights continuous and learnable. We introduce a meta-loss that jointly optimizes for attack effectiveness and perceptual similarity to human vision. This allows the weights to be adjusted through gradient-based optimization, effectively converting norm-constrained base attacks into perceptually aligned attacks—without needing to modify their core mechanisms.
Regarding Observation 2, we mitigate the risk of poor local optima by introducing a multi-stage chaining process. In each stage, we recombine base attacks using newly learned weights and feed the output adversarial example from the previous stage into the next as input. This design allows the attack to iteratively refine the adversarial perturbation. Prior work on transfer-based attacks has shown the benefits of chaining attacks across stages Suya et al. (2020), but typically only use two stages and assign different attacks to each. In contrast, our method allows arbitrary depth and continuously optimized attack combinations across all stages.
4.2 Design Details
Next, we introduce our attack design more formally. We assume a total of $M$ base attacks and $N$ stages (or attack cells), both of which are hyperparameters that the attacker can freely choose to optimize attack performance. The set of base attacks can also be easily expanded as new attack methods become available in the future.
In a given stage $j$, we denote the learnable weight for each base attack as $\alpha_{j,i}$, where $i$ indexes the base attack. The output of the $j$-th attack cell is then defined as:
$\mathcal{A}_{j}(\boldsymbol{x}^{j-1}_{a},y)=\sum_{i=1}^{M}\frac{\exp(\alpha_{j,i})}{\sum_{k=1}^{M}\exp(\alpha_{j,k})}\cdot\boldsymbol{x}_{a}^{j,i}$ (3)
where $\boldsymbol{x}_{a}^{j,i}$ denotes the adversarial example generated by the $i$-th base attack at stage $j$. The softmax operation ensures that the final combination remains within the original $\ell_{p}$-norm constraints, as we aim to preserve the structure of the base attacks without modifying them directly. This linear formulation provides a simple and extensible mechanism for combining base attacks. Future work may explore more advanced combination strategies beyond weighted averaging, or treat individual hyperparameters of base attacks, which are fixed in this work, as additional learnable variables.
After the adversarial example generation in each stage, we apply multi-stage chaining as follows: $\boldsymbol{x}_{a}=\mathcal{A}_{N}(\boldsymbol{x}_{a}^{N-1},y)\circ\cdots\circ\mathcal{A}_{1}(\boldsymbol{x}^{0},y)$ where the composition operator $\circ$ indicates the output of each stage is passed as input to the next. This enables iterative refinement of the adversarial perturbation over multiple stages. While we adopt a simple linear chaining strategy here, more sophisticated chaining mechanisms (e.g., cyclic or attention) are possible and left as future work.
Finally, we define the meta-loss that guides the optimization of the attack weights: $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{ssim}}\cdot(1-\text{SSIM}(\boldsymbol{x},\boldsymbol{x}_{a}))$ The first term encourages attack effectiveness by maximizing the confidence drop in the ground-truth class $y$ for the adversarial example $\boldsymbol{x}_{a}$. The second term penalizes perceptual distortion, measured using the Structural Similarity Index Measure (SSIM), which ranges from 0 to 1, with higher values indicating greater similarity. In practice, we observe that substituting SSIM with other perceptual similarity metrics yields comparable results. Our full attack algorithm is summarized in Algorithm 1 and a visual demonstration is also shown in Figure 2.
Algorithm 1 Search for Optimal Attack Sequence $\mathcal{A}_{\Theta}$
1:Randomly Initialize $\boldsymbol{\alpha}$
[…]
3:for $\text{epoch}=1$ to $T$ do
4:  Sample a mini-batch $\{(\boldsymbol{x},y)\}\subset\mathcal{D}$
5:  Generate adversarial examples: $\boldsymbol{x}_{a}\leftarrow\mathcal{A}(\boldsymbol{x},y)$
6:  Feed to target model: $f(\boldsymbol{x}_{a})$
7:  Compute total loss $\mathcal{L}_{\text{total}}$
8:  Update parameters: $\boldsymbol{\alpha}\leftarrow\boldsymbol{\alpha}-\eta\,\nabla_{\boldsymbol{\alpha}}\mathcal{L}_{\text{total}}(\boldsymbol{x}_{a},y)$
9:end for
10:Return attack sequence $\mathcal{A}_{\Theta}$ that minimizes $\mathcal{L}_{\text{total}}$
Table 1: Performance comparison of adversarial attacks across different datasets and RobustBench models. Base is the ASR without applying any post-processing defenses; Avg. is the average of all the ASR values. SSIM values are scaled to [0,100].
Dataset Target Model Attack ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
Base JPEG TVM Ensemble NRP Avg.
CIFAR-100 Cui et al. (2024) DI-FGSM 65.20 65.49 65.35 65.55 57.03 63.72 92.08 0.0168 54.93
CW 100.00 87.70 96.68 78.61 57.71 84.14 94.38 0.0153 32.49
PI-FGSM++ 64.41 65.78 64.32 65.64 50.86 62.20 91.12 0.0199 58.94
AutoAttack 82.56 81.33 79.51 80.55 59.88 76.77 92.09 0.0167 41.68
DiffPGD 76.20 76.00 75.80 75.60 75.40 75.80 91.15 0.0165 43.50
DiffAttack 77.50 77.20 77.70 77.40 76.40 77.24 91.16 0.0169 42.30
AdvAD 82.20 82.80 81.10 81.70 67.90 79.14 83.18 0.0288 45.67
[…]
Wang et al. (2023) DI-FGSM 66.46 67.01 66.07 66.72 59.86 65.22 91.96 0.0169 55.24
CW 100.00 85.74 95.31 76.95 59.86 83.57 94.55 0.0161 32.19
PI-FGSM++ 66.17 67.21 65.29 66.76 52.30 63.55 91.14 0.0199 57.88
AutoAttack 83.50 82.91 81.05 81.64 62.60 78.34 91.25 0.0189 42.19
DiffPGD 79.10 76.10 74.20 73.80 73.40 75.32 91.50 0.0161 51.80
DiffAttack 76.90 76.60 76.40 76.80 76.30 76.53 90.90 0.0176 42.20
AdvAD 79.90 79.60 79.90 79.50 67.20 77.42 83.07 0.0293 47.29
[…]
(Addepalli et al.) (2022) DI-FGSM 72.09 72.13 71.84 71.74 63.42 70.24 92.05 0.0156 56.69
CW 100.00 87.21 96.48 77.93 60.74 84.47 95.56 0.0113 26.94
PI-FGSM++ 71.50 71.93 70.86 71.15 56.89 68.47 91.14 0.0192 59.94
AutoAttack 83.77 82.70 81.02 80.68 61.33 77.90 93.03 0.0147 39.59
DiffPGD 82.50 82.20 81.80 81.60 80.90 81.80 91.05 0.0168 42.80
DiffAttack 81.60 81.60 81.40 81.40 81.50 81.35 91.26 0.0163 42.15
AdvAD 84.50 84.00 83.70 84.30 70.50 81.60 83.92 0.0291 45.63
[…]
CIFAR-10 (Bartoldson et al.) (2024) DI-FGSM 41.39 42.46 42.27 42.75 35.70 40.91 90.88 0.0200 59.07
CW 97.27 61.62 96.09 60.35 58.59 74.78 95.25 0.0206 55.25
PI-FGSM++ 45.80 48.46 46.72 48.89 23.24 42.62 88.74 0.0278 66.91
AutoAttack 73.24 71.88 72.75 71.19 39.36 65.68 87.51 0.0264 53.05
DiffPGD 53.20 52.40 52.60 52.40 52.80 52.68 91.60 0.0165 54.20
DiffAttack 73.60 73.30 73.60 73.50 72.90 73.38 82.90 0.0369 51.60
AdvAD 61.30 60.70 60.60 61.20 41.10 56.98 83.50 0.0269 49.37
[…]
Cui et al. (2024) DI-FGSM 51.64 52.71 52.36 53.01 45.23 50.99 90.51 0.0206 62.20
CW 99.80 67.68 95.51 65.33 65.33 78.73 95.39 0.0206 51.38
PI-FGSM++ 55.76 57.09 56.05 56.84 31.07 51.36 88.71 0.0272 68.81
AutoAttack 78.42 77.34 78.22 76.46 44.34 70.96 86.88 0.0273 56.71
DiffPGD 76.20 75.80 76.50 76.10 75.40 76.00 79.80 0.0350 60.50
DiffAttack 82.80 82.50 83.10 82.70 82.20 82.66 81.36 0.0404 58.24
AdvAD 66.80 67.20 66.60 66.20 45.50 62.46 83.11 0.0265 50.95
[…]
Wang et al. (2023) DI-FGSM 51.64 52.13 52.09 51.84 44.57 50.45 90.45 0.0203 63.21
CW 98.90 70.21 94.82 66.11 51.76 76.36 95.25 0.0208 52.86
PI-FGSM++ 55.29 57.11 54.61 56.27 29.61 50.58 88.72 0.0265 70.71
AutoAttack 77.93 77.54 77.73 76.27 43.16 70.53 86.83 0.0276 56.75
DiffPGD 75.80 75.40 76.10 75.60 75.10 75.60 79.50 0.0345 60.80
DiffAttack 82.50 82.40 82.50 82.40 82.10 82.38 80.12 0.0429 57.22
AdvAD 67.40 68.40 66.40 67.20 44.10 62.70 82.64 0.0271 52.78
[…]
ImageNet-1k Salman et al. (2020) DI-FGSM 83.90 83.60 83.00 82.60 75.40 81.70 91.42 0.0849 49.39
CW 98.50 88.40 98.20 76.00 63.70 85.02 97.75 0.0263 6.77
PI-FGSM++ 87.10 86.80 85.60 85.10 74.50 83.82 90.54 0.0960 49.57
AutoAttack 94.90 94.30 91.30 90.30 72.60 91.56 92.90 0.0757 21.78
DiffPGD 76.20 69.40 74.80 68.50 72.30 72.24 87.50 0.0892 45.60
DiffAttack 92.50 86.80 88.90 87.90 89.80 89.18 89.17 0.0763 40.78
AdvAD 89.20 87.90 87.70 87.90 86.80 87.90 71.30 0.1949 50.14
[…]
5 Experiments and Results
Next we discuss our experimental setup and evaluations. All the reported results are obtained using NVIDIA A100 GPU.
5.1 Experimental Setup
5.1.1 Datasets and Models
Our experiments are conducted on three commonly used benchmark datasets of different scales: CIFAR-10 Krizhevsky et al. (2014), CIFAR-100 Krizhevsky (2009), and ImageNet Deng et al. (2009). For each dataset, we randomly sample 1,000 test images that are correctly classified by the victim model. Since standard models (without any defenses) are easily broken by existing attacks, we focus our evaluation on robust models curated by RobustBench Croce et al. (2020), a popular benchmark for adversarial robustness.
5.1.2 Post-Processing Defenses
To further test the limits of DAASH and other SOTA attacks in terms of generating potent adversarial examples, we also experiment with post-processing defenses (on top of using RobustBench models). The post-processing defenses include JPEG compression Guo et al. (2017), Total Variation Minimization (TVM) Wang et al. (2018), Neural Representation Purifier (NRP) Naseer et al. (2020), and an ensemble of four methods: bit-depth reduction Xu (2018), JPEG, TVM, and Non-Local Means Denoising (NLM) Buades et al. (2011).
5.1.3 Evaluation Metrics
We use attack success rate (ASR) to measure attack potency, and for determining human visual imperceptibility, we use: SSIM, LPIPS, and FID.
Figure 3: Comparison of DAASH performance before and after training, illustrating ASR and SSIM across different defense settings. “Trained DAASH" significantly outperforms the randomly weighted version (“Random DAASH").
5.1.4 Base Attacks and DAASH Training Details
We include 10 diverse base attacks in our attack pool: FGSM Goodfellow et al. (2015b), PGD Madry et al. (2018b), CW Carlini and Wagner (2017), FAB Croce and Hein (2020a), TI-FGSM Dong (2019), NI-FGSM Lin et al. (2019), MI-FGSM Dong et al. (2017), DI-FGSM Xie et al. (2018), BIM Kurakin et al. (2016), and PI-FGSM++ Gao et al. (2020). Each attack offers unique advantages and covers a broad range of perturbation strategies. In addition, we include a None operation at each stage, which simply returns the input image, allowing the model to flexibly decide whether to skip processing in a given stage.
For the meta-loss, we set $\lambda_{\text{asr}}=1.3$ and $\lambda_{\text{ssim}}=1.0$, which work well across datasets. We run Algorithm 1 for $T=100$ epochs using the Adam optimizer with a learning rate of 0.001. After training, the learned weights for combining base attacks in each stage are used to generate the final adversarial example.
Figure 4: DAASH training progression over epochs.
5.2 Effectiveness of DAASH against Robust Models
Table 1 reports the attack success rate (ASR) and perceptual similarity scores of DAASH and seven representative state-of-the-art attacks evaluated against seven recent adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet. As shown in the table, DAASH consistently outperforms all baseline attacks in terms of ASR under both the Base (no post-processing) and post-processing defense conditions, across all datasets. This includes outperforming recent attacks explicitly designed to go beyond $\ell_{p}$-norm constraints. For example, on CIFAR-100, the best-performing baseline is CW with an average ASR of $84\%$, while the most recent perceptual-aligned attack, AdvAD, achieves only $79\%$. In contrast, DAASH achieves an ASR close to $100\%$. Moreover, DAASH yields better imperceptibility scores compared to these dedicated perceptual attacks. Similar trends are observed on CIFAR-10 and ImageNet: DAASH improves over the best baseline by at least $15\%$ ASR on CIFAR-10 and $5\%$ on ImageNet, on average. Its perceptual scores are also consistently better than other attacks specifically designed for perceptual alignment.
Table 2: Effectiveness of DAASH with varying number of stages for the robust CIFAR-100 model – Cui et al. (2024).
Num. of ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
Stages Base JPEG TVM Ensemble NRP Avg.
1 72.85 34.08 68.46 37.70 46.29 51.88 98.48 0.0027 16.19
2 100.00 99.90 100.00 98.54 99.02 99.49 94.59 0.0115 43.28
3 100.00 99.90 100.00 99.32 99.61 99.77 94.43 0.0139 41.08
4 100.00 100.00 100.00 99.61 99.71 99.86 93.22 0.0190 47.56
5 100.00 100.00 100.00 100.00 99.90 99.98 92.92 0.0209 48.25
Table 3: Average attack success rate (ASR↑) and structural similarity (SSIM↑) for cross-model transferability.
Trained Tested CIFAR-100 CIFAR-10 ImageNet-1K
(Cui et al.) (2024) (Wang et al.) (2023) (Addepalli et al.) (2022) (Bartoldson et al.) (2024) (Cui et al.) (2024) (Wang et al.) (2023) (Salman et al.) (2020)
Dataset ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM
[…]
Effectiveness of CW. An interesting finding is that the CW attack, originally developed for $\ell_{2}$-norm constrained settings, also achieves competitive ASR and perceptual similarity when its hyperparameters are carefully tuned (we tested multiple configurations and report the best result). In some cases, CW was also able to outperform DiffPGD, AdvAD, and DiffAttack. This suggests that some classical attacks, when properly tuned, can potentially still be highly effective in terms of both ASR and imperceptibility.
Although DAASH outperforms CW in terms of ASR, its imperceptibility scores are sometimes slightly lower. Intuitively, since CW is one of the base attacks in DAASH, it could be expected to receive higher attention weights during optimization. However, this is not always the case. The reason lies in the mismatch between the objectives: CW uses a hinge-like loss that stops contributing gradients once misclassification is achieved, whereas our meta-loss continues to penalize the confidence of the true class beyond the decision boundary. As a result, after initial misclassification, CW receives lower weights in DAASH, which shifts focus to other attacks that still produce informative gradients. This continued “pushing” of adversarial examples beyond the boundary enhances robustness of DAASH against additional defenses such as post-processing but comes at the cost of slightly reduced perceptual similarity.
5.3 Ablation Studies on DAASH
Our DAASH framework consists of two key components: (1) learnable weights for each base attack within every attack cell, and (2) chained attack cells to better escape poor local optima. In this section, we conduct ablation studies to demonstrate the importance of each component.
Importance of Learnable Weights. To evaluate the benefit of using learnable weights, we compare the performance of DAASH against a variant where the weights for base attacks are initialized randomly and left untrained. Specifically, the random weights are sampled uniformly in the range $[0,1]$ and used directly without optimization. We repeat this random setup 100 times and report the average results.
Figure 3 shows that using trained weights significantly improves both ASR and imperceptibility (measured by SSIM) across all types of robust models. Furthermore, Figure 4 illustrates the ASR and SSIM trends as training progresses over epochs. Both metrics exhibit clear improvement—ASR increases across all settings, and SSIM improves almost monotonically—highlighting the effectiveness of our meta-learning process. These results confirm that the learnable weighting mechanism in DAASH plays a crucial role.
Importance of Multiple Stages. To assess the value of chaining multiple stages, we evaluate DAASH with varying numbers of stages. Table 2 reports the ASR for different stage configurations. The results show a substantial ASR jump when increasing from one to two stages (from 51.88% to 99.49%), suggesting that two stages are sufficient to escape bad local optima and refine the adversarial example generation. Adding more stages yields marginal improvements, indicating diminishing returns beyond two stages.
Effectiveness against Unseen Defenses. We also examine whether the learned weights in DAASH are transferable across models and datasets. Table 3 shows that weights trained on one model or dataset can be applied to unseen models or datasets while still achieving near-optimal results. The variation in average ASR remains within 3%, and SSIM varies by less than 2. These results demonstrate that DAASH can generalize well across different settings/models.
6 Conclusion
In this work, we introduced DAASH, a novel differentiable framework for learning optimal compositions of adversarial attacks in a white-box setting. Unlike traditional handcrafted methods that heuristically combine attacks, DAASH leverages gradient-based optimization to learn soft weightings over multiple attack modules across multiple stages. This process is guided by a meta-loss that jointly captures attack effectiveness and perceptual alignment with human vision.
Our flexible framework enables the transformation of existing norm-constrained attacks into stronger attacks that also produce perceptually aligned adversarial examples. Empirical results demonstrate that DAASH  significantly outperforms state-of-the-art perceptually aligned attacks in both attack success rate and imperceptibility scores. DAASH  generalizes well to unseen defenses, establishing itself as a strong and adaptive baseline for future robustness evaluation.
Overall, our results underscore the potential of differentiable search in the adversarial example literature and open up new directions for developing more adaptive and powerful attack strategies. Future work may explore extending DAASH  to black-box or query-limited scenarios, as well as integrating it with generative models to further enhance imperceptibility and increase sample diversity.

### v2

@@
-DAASH: A Meta-Attack Framework for Synthesizing Effective and Stealthy Adversarial Examples
+DASH: A Meta-Attack Framework for Synthesizing Effective and Stealthy Adversarial Examples
 Abstract
-Numerous techniques have been proposed for generating adversarial examples in white-box settings under strict $\ell_{p}$-norm constraints.
+Numerous techniques have been proposed for generating adversarial examples under strict $\ell_{p}$-norm constraints.
 However, such norm-bounded examples often fail to align well with human perception, and only recently have a few methods begun specifically exploring perceptually aligned adversarial examples.
 Moreover, it remains unclear whether insights from $\ell_{p}$-constrained attacks can be effectively leveraged to improve perceptual efficacy.
-In this paper, we introduce DAASH, a fully differentiable meta-attack framework that generates effective and perceptually aligned adversarial examples by strategically composing existing $\ell_{p}$-based attack methods.
-DAASH operates in a multi-stage fashion: at each stage, it aggregates candidate adversarial examples from multiple base attacks using learned, adaptive weights and propagates the result to the next stage.
+In this paper, we introduce DASH, a differentiable meta-attack framework that generates effective and perceptually aligned adversarial examples by strategically composing existing $\ell_{p}$-based attack methods.
+DASH operates in a multi-stage fashion: at each stage, it aggregates candidate adversarial examples from multiple base attacks using learned, adaptive weights and propagates the result to the next stage.
 A novel meta-loss function guides this process by jointly minimizing misclassification loss and perceptual distortion, enabling the framework to dynamically modulate the contribution of each base attack throughout the stages.
-We evaluate DAASH on adversarially trained models across CIFAR-10, CIFAR-100, and ImageNet.
-Despite relying solely on $\ell_{p}$-constrained based methods, DAASH significantly outperforms state-of-the-art perceptual attacks such as AdvAD—achieving higher attack success rates (e.g., 20.63% improvement) and superior visual quality, as measured by SSIM, LPIPS, and FID (improvements of $\approx 11$, 0.015, and 5.7, respectively).
-Furthermore, DAASH generalizes well to unseen defenses, making it a practical and strong baseline for evaluating robustness without requiring handcrafted adaptive attacks for each new defense.
+We evaluate DASH on adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet while considering visual perception metrics (e.g.
+SSIM, FID, LPIPS) in the perturbation budget (instead of $\ell_{p}$-norm).
+Despite relying solely on $\ell_{p}$-constrained based methods, DASH significantly outperforms state-of-the-art perceptual attacks such as AdvAD, achieving higher attack success rates (e.g., 20.63% improvement) and superior visual quality, as measured by SSIM, LPIPS, and FID (improvements of $\approx 11$, 0.015, and 5.7, respectively).
+DASH generalizes well to unseen defenses and different white-box/black-box scenarios, making it a practical and strong baseline for evaluating robustness.
 1 Introduction
-Adversarial examples against deep neural networks (DNNs), which add imperceptible perturbations to clean images to induce misclassification (targeted/untargeted) Goodfellow et al.
-(2015a), have attracted significant attention from the research community.
-Various attack methods, including FGSM Goodfellow et al.
-(2015b), PGD Madry et al.
-(2018b), and CW Carlini and Wagner (2017), as well as adaptive evaluation frameworks such as AutoAttack Croce et al.
-(2020), have been developed to benchmark both defended and undefended models.
+Adversarial examples against deep neural networks (DNNs), which add imperceptible perturbations to clean images to induce misclassification (targeted/untargeted) [goodfellow2014explaining], have attracted significant attention from the research community.
+Various attack methods, including FGSM [goodfellow2015explaining], PGD [madry2018towards], and CW [carlini2017towards], as well as adaptive evaluation frameworks such as AutoAttack [croce2020robustbench], have been developed to benchmark both robust and undefended models.
 However, the majority of these approaches prioritize imperceptibility by constraining perturbations within a small $\ell_{p}$-norm ball, with $\ell_{2}$ and $\ell_{\infty}$ norms being the most common ones.
-Yet, minimizing $\ell_{p}$-norm does not always yield high perceptual similarity to the human visual system Zhang et al.
-(2018); Laidlaw and Feizi (2020); Sharif et al.
-(2018), despite imperceptibility being a fundamental design goal of stealthy adversarial examples.
-This disconnect has motivated the adoption of perceptually-aligned similarity metrics such as Structural Similarity Index (SSIM) Wang et al.
-(2004), Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al.
-(2018), and Fréchet Inception Distance (FID) Heusel et al.
-(2017) to better assess visual quality Li et al.
-(2024); Chen et al.
-(2024b).
-However, attacks optimized under these perceptual metrics often suffer from reduced success rates—particularly against robust, defended models (see Table 1)—highlighting a critical trade-off and leaving significant room for improvement.
-Designing effective attacks against a wide range of models, especially hardened ones, remains non-trivial and often requires deep, manual effort—an observation consistently echoed in the literature on $\ell_{p}$-norm attacks Carlini and Wagner (2017); Athalye et al.
-(2018); Croce and Hein (2020b).
+Yet, minimizing $\ell_{p}$-norm does not always yield high perceptual similarity to the human visual system [zhang2018unreasonable, laidlaw2020perceptual, sharif2018suitability], despite imperceptibility being a fundamental design goal of stealthy adversarial examples.
+This disconnect has motivated the adoption of perceptually-aligned similarity metrics such as Structural Similarity Index (SSIM) [wang2004image], Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable], and Fréchet Inception Distance (FID) [heusel2017gans] to better assess visual quality [li2024advad, chen2024diffusion].
+However, attacks optimized under these perceptual metrics often suffer from reduced success rates; particularly against robust, defended models (see Table 1), highlighting a critical trade-off and leaving significant room for improvement.
+Designing effective attacks against a wide range of models, especially hardened ones, remains non-trivial and often requires deep, manual effort, an observation consistently echoed in the literature on $\ell_{p}$-norm attacks [carlini2017towards, athalye2018obfuscated, croce2020reliable].
 This complexity intensifies when attempting to simultaneously preserve high perceptual fidelity, thus underscoring the need for systematic and automated strategies that can generalize across models while aligning better with human vision.
-Figure 1: Comparative visualization of adversarial perturbations across multiple attack methods (including DAASH).
+Figure 1: Comparative visualization of adversarial perturbations across multiple attack methods (including DASH).
 Given the extensive body of work on $\ell_{p}$-bounded attacks, including those targeting robust models, it is natural to ask:
 Can we leverage the strengths of these existing attacks to compose stronger adversarial examples that are also more perceptually aligned?
-Our key insight is that although individual $\ell_{p}$-norms (e.g., $\ell_{2}$, $\ell_{\infty}$) do not fully capture human perception, they each correlate with different perceptual characteristics (e.g., contrast, texture, edge disruption) Sharif et al.
-(2018); Laidlaw and Feizi (2020).
+Our key insight is that although individual $\ell_{p}$-norms (e.g., $\ell_{2}$, $\ell_{\infty}$) do not fully capture human perception, they each correlate with different perceptual characteristics (e.g., contrast, texture, edge disruption) [sharif2018suitability, laidlaw2020perceptual].
 Therefore, by combining attack strategies optimized under different norms in a principled manner, we may achieve not only improved attack success but also better perceptual quality, particularly when targeting defended models where individual attack methods may fall short.
-Specifically, we propose DAASH (Differential Adversarial Attack SearcH)—a fully differentiable, multi-stage framework that learns to compose multiple $\ell_{p}$-norm–constrained base attacks using soft attention.
-DAASH operates in two stages: (1) it computes a soft attention distribution over adversarial examples generated from individual candidate attacks (e.g., PGD, CW) to weight and compose the final adversarial example; (2) it iteratively refines the perturbations across multiple stages by feeding the output from one stage as the input to the next.
+Specifically, we propose DASH (Differential Attack SearcH), a fully differentiable, multi-stage framework that learns to compose multiple $\ell_{p}$-norm–constrained base attacks using soft attention.
+DASH operates in two stages: (1) it computes a soft attention distribution over adversarial examples generated from individual candidate attacks (e.g., PGD, CW) to weight and compose the final adversarial example; (2) it iteratively refines the perturbations across multiple stages by feeding the output from one stage as the input to the next.
 This iterative chaining is designed to escape poor local minima inherent to the non-convex landscape of adversarial optimization.
@@
 This allows the framework to dynamically adjust the influence of each candidate attack based on the target model.
-Unlike existing methods that are rigid in balance between imperceptibility (from human perception) and attack success, DAASH learns to trade off these objectives adaptively across different models and datasets.
-Empirically, DAASH achieves a significant improvement in both attack success rate (ASR) and perceptual fidelity compared to the current state-of-the-art.
-On robust models trained on CIFAR-100 such as Cui et al.
-(2024), DAASH reaches an average ASR of 99.97% with an SSIM score of 94.43, significantly outperforming state-of-the-art perceptual attacks such as AdvAD Li et al.
-(2024) (ASR: 79.14%, SSIM: 83.18) and DiffAttack Chen et al.
-(2024b) (ASR: 77.24%, SSIM: 91.16).
+Unlike existing methods that are rigid in balance between imperceptibility (from human perception) and attack success, DASH learns to trade off these objectives adaptively across different models and datasets.
+Empirically, DASH achieves a significant improvement in both attack success rate (ASR) and perceptual fidelity compared to the current state-of-the-art.
+On robust models trained on CIFAR-100 such as [cui2024decoupled], DASH reaches an average ASR of 99.77% with an SSIM score of 94.43, significantly outperforming state-of-the-art perceptual attacks such as AdvAD [li2024advad] (ASR: 79.14%, SSIM: 83.18) and DiffAttack [chen2024diffusion] (ASR: 77.24%, SSIM: 91.16).
 Similar improvements are observed in other experimental settings.
-Fig. 1 presents qualitative comparisons, further illustrating the superiority of DAASH in generating perceptually plausible yet highly effective adversarial examples.
+DASH also transfers well across different models, which demonstrates its effectiveness in black-box settings.
+Fig. 1 presents qualitative comparisons, further illustrating the superiority of DASH in generating perceptually plausible yet highly effective adversarial examples.
 Our contributions are summarized as follows:
-  • We propose DAASH, a fully differentiable, multi-stage framework that learns to combine $\ell_{p}$-norm–constrained base attacks using soft attention.
-  • We introduce a novel meta-loss that decouples attack effectiveness and human perceptual quality from the constraints of individual norm-based attacks, enabling better identification of attack strengths across models.
-  • We validate DAASH across multiple datasets (CIFAR-10, CIFAR-100, ImageNet), seven robust models, and four post-processing defenses demonstrating substantial improvements over prior perceptual attacks.
+  • Propose DASH, a differentiable, multi-stage meta-framework that learns to combine multiple base attacks using soft attention.
+DASH decouples the objectives of attack effectiveness and stealth (e.g., SSIM), allowing for custom optimization of each using a novel meta-loss.
+  • Validate DASH across multiple datasets (CIFAR-10, CIFAR-100, ImageNet), seven robust models, and four post-processing defenses demonstrating substantial improvements over prior perceptual attacks.
+  • Demonstrate the effectiveness of DASH in black-box settings through rigorous transferability analysis.
 2 Background and Threat Model
@@
 In the equation above, $d(\cdot,\cdot)$ represents a distance metric and $\epsilon$ defines the perturbation budget.
-$d(\cdot,\cdot)$ measures the imperceptibility of the perturbation, and the commonly used metrics include $\ell_{2}$ norm Carlini and Wagner (2017) as well as the $\ell_{\infty}$-norm Carlini and Wagner (2017).
-These $\ell_{p}$-norm constraints may not always align with human perception, and some alternative metrics such as the Structural Similarity Index (SSIM) Wang et al.
-(2004) is used.
+$d(\cdot,\cdot)$ measures the imperceptibility of the perturbation, and the commonly used metrics include $\ell_{2}$ norm [carlini2017towards] as well as the $\ell_{\infty}$-norm [carlini2017towards].
+These $\ell_{p}$-norm constraints may not always align with human perception, and some alternative metrics such as the Structural Similarity Index (SSIM) [wang2004image] is used.
 SSIM is shown to better capture the luminance, contrast, and structural similarities between two images $\boldsymbol{x}_{1},\boldsymbol{x}_{2}$ and is computed as below:
@@
 The constants $c_{1}$ and $c_{2}$ are used to stabilize the division when the denominator is close to zero.
-Besides SSIM, other perceptual similarity metrics that better align with human perception include the Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al.
-(2018) and the Fréchet Inception Distance (FID) Heusel et al.
-(2017).
-In this paper, we use SSIM as the primary metric to optimize for imperceptibility, but we find that using LPIPS and FID yields similar results in practice.
+Besides SSIM, other perceptual similarity metrics that better align with human perception include the Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable] and the Fréchet Inception Distance (FID) [heusel2017gans].
+In this paper, we use SSIM as the perturbation constraint, it’s the primary metric to optimize for imperceptibility, but we find that using LPIPS also yields similar results in practice.
+Using FID in the optimize process is possible but not ideal in our current experimental setup (small batch FID calculation can be noisy).
 Threat model.
-In this paper, we focus on the white-box setting, where the attacker is assumed to have complete access to $f(\cdot)$, including the model architecture and parameters.
+In this paper, we focus on both white-box and black-box settings.
+For white-box, the attacker is assumed to have complete access to $f(\cdot)$, including the model architecture and parameters.
 In this setting, untargeted adversarial examples, as defined in Eqn.
 (1), can be generated via backpropagation using well-defined loss functions with respect to the input $\boldsymbol{x}$.
-The white-box assumption is widely adopted in the extensive literature Carlini and Wagner (2017); Madry et al.
-(2018a); Croce and Hein (2020b).
+White-box assumption is the default for robustness analysis (e.g CW [carlini2017towards], AutoAttack [croce2020reliable]) and is also considered in recent (closely related) perception-aligned attacks such as DiffPGD [xue2023diffpgd] (NeurIPS’23) and AdvAD [li2024advad] (NeurIPS’24).
+For black-box, It is assumed that the attacker does not have any access to the target model.
+We consider generating the adversarial examples using a surrogate model and then it’s effect is measured on the target model.
+Note that transfer attacks are under even stricter threat model than query-based black-box attacks.
 3 Related Work
-Classical approaches for generating adversarial examples include FGSM Goodfellow et al.
-(2015a), I-FGSM Kurakin et al.
-(2016), and PGD Madry et al.
-(2018a), all of which strictly enforce a perturbation budget $\epsilon$ measured in the $\ell_{\infty}$-norm.
-In contrast, the CW attack Carlini and Wagner (2017) softly enforces an $\ell_{2}$-norm perturbation constraint by incorporating the perturbation magnitude directly into the optimization objective.
-Closer to our approach are AutoAttack Croce and Hein (2020b) and Composite Attacks Mao et al.
-(2021).
+Classical approaches for generating adversarial examples include FGSM [goodfellow2014explaining], I-FGSM [kurakin2016adversarial], and PGD [madry2017towards], all of which strictly enforce a perturbation budget $\epsilon$ measured in the $\ell_{\infty}$-norm.
+In contrast, the CW attack [carlini2017towards] softly enforces an $\ell_{2}$-norm perturbation constraint by incorporating the perturbation magnitude directly into the optimization objective.
+Closer to our approach are AutoAttack [croce2020reliable] and Composite Attacks [mao2021composite].
 AutoAttack employs an ensemble of a few selected attack methods, while composite attacks combine multiple attacks using genetic algorithms.
 However, these approaches still operate under conventional $\ell_{p}$-norm constraints and rely on simplistic or heuristic strategies for combining base attacks.
-In contrast, our proposed method DAASH introduces an optimization-based meta-framework to adaptively compose base attacks for improved attack effectiveness and imperceptibility for human eyes.
+In contrast, our proposed method DASH introduces an optimization-based meta-framework to adaptively compose base attacks for improved attack effectiveness and imperceptibility for human eyes.
 Recognizing the limitations of norm-constrained attacks in modeling human perception, recent works have shifted toward leveraging perceptual similarity metrics to guide adversarial example generation.
-For instance, SSAH Luo et al.
-(2022) operates in the frequency domain to introduce perturbations in less perceptible high-frequency components, while PerC-AL Zhao et al.
-(2020) optimizes perceptual color distance, though it relies on manual hyperparameter tuning.
+For instance, SSAH [luo2022frequency] operates in the frequency domain to introduce perturbations in less perceptible high-frequency components, while PerC-AL [zhao2020towards] optimizes perceptual color distance, though it relies on manual hyperparameter tuning.
 More recent approaches exploit diffusion models by injecting perturbations into their embedding spaces, thereby influencing the denoising trajectory to produce adversarial outputs.
-Representative examples include DiffAttack Chen et al.
-(2024a), AdvAD Li et al.
-(2024), and DiffPGD Xue et al.
-(2023).
+Representative examples include DiffAttack [chen2023diffusion], AdvAD [li2024advad], and DiffPGD [xue2023diffusion].
 Despite their innovations, these perceptual attacks are often designed as standalone strategies with limited adaptability, making them less effective across diverse target models.
 As demonstrated in Table 1, such methods frequently underperform in both attack success and imperceptibility.
-By contrast, DAASH builds on the extensive literature of norm-constrained attacks—known for their compatibility with a wide range of target models—and combines them in a unified framework that enhances the effectiveness and stealth of the resulting adversarial examples.
-4 DAASH Attack Design
-Next, we describe the high-level motivation behind the design of DAASH and then introduce the details more formally.
-Figure 2: An overview of the proposed Differentiable Adversarial Attack Search (DAASH) methodology.
+By contrast, DASH builds on the extensive literature of norm-constrained attacks, known for their compatibility with a wide range of target models and combines them in a unified framework that enhances the effectiveness and stealth of the resulting adversarial examples.
+4 DASH Attack Design
+Next, we describe the high-level motivation behind the design of DASH and then introduce the details more formally.
+Figure 2: An overview of the proposed Differentiable Attack Search (DASH) methodology.
 4.1 Design motivation
-Our attack design is motivated by two key observations: (1) Norm-constrained attacks sometimes exhibit complementary strengths and weaknesses Croce and Hein (2020b).
-So, combining them can yield better performance, especially in under-explored settings of generating adversarial examples beyond $\ell_{p}$-norm constraints; (2) Adversarial example generation is a non-convex optimization process and is prone to getting stuck in poor local optima Madry et al.
-(2018a).
+Our attack design is motivated by two key observations: (1) Norm-constrained attacks sometimes exhibit complementary strengths and weaknesses [croce2020reliable].
+So, combining them can yield better performance, especially in under-explored settings of generating adversarial examples beyond $\ell_{p}$-norm constraints; (2) Adversarial example generation is a non-convex optimization process and is prone to getting stuck in poor local optima [madry2017towards].
 Regarding Observation 1, the most natural way to combine different attacks is to take a weighted average of the adversarial examples generated by each individual attack.
-While a linear combination is simple and practical, it introduces two key challenges: (a) How to determine optimal weights for each base attack, and (b) How to ensure that the resulting adversarial example better aligns with human perception rather than just staying within the $\ell_{p}$ constraints.
+While a linear combination is practical, it introduces two key challenges: (a) How to determine optimal weights for each base attack, and (b) How to ensure that the resulting adversarial example better aligns with human perception rather than just staying within the $\ell_{p}$ constraints.
 To address these, we make the combination weights continuous and learnable.
 We introduce a meta-loss that jointly optimizes for attack effectiveness and perceptual similarity to human vision.
-This allows the weights to be adjusted through gradient-based optimization, effectively converting norm-constrained base attacks into perceptually aligned attacks—without needing to modify their core mechanisms.
+This allows the weights to be adjusted through gradient-based optimization, effectively converting norm-constrained base attacks into perceptually aligned attacks, without needing to modify their core mechanisms.
 Regarding Observation 2, we mitigate the risk of poor local optima by introducing a multi-stage chaining process.
@@
 This design allows the attack to iteratively refine the adversarial perturbation.
-Prior work on transfer-based attacks has shown the benefits of chaining attacks across stages Suya et al.
-(2020), but typically only use two stages and assign different attacks to each.
+Prior work on transfer-based attacks has shown the benefits of chaining attacks across stages [suya2020hybrid], but typically only use two stages and assign different attacks to each.
 In contrast, our method allows arbitrary depth and continuously optimized attack combinations across all stages.
+Algorithm 1 Search for Optimal Attack Sequence $\mathcal{A}_{\Theta}$
+1:Randomly Initialize $\boldsymbol{\alpha}$
+[…]
+3:for $\text{epoch}=1$ to $T$ do
+4:  Sample a mini‑batch $\{(\boldsymbol{x},y)\}\subset\mathcal{D}$
+5:  Generate adversarial examples: $\boldsymbol{x}_{a}\leftarrow\mathcal{A}(\boldsymbol{x},y)$
+6:  Feed to target model: $f(\boldsymbol{x}_{a})$
+7:  Compute total loss $\mathcal{L}_{\text{total}}$
+8:  Update parameters: $\boldsymbol{\alpha}\leftarrow\boldsymbol{\alpha}-\eta\,\nabla_{\boldsymbol{\alpha}}\mathcal{L}_{\text{total}}(\boldsymbol{x}_{a},y)$
+9:end for
+10:Return attack sequence $\mathcal{A}_{\Theta}$ that minimizes $\mathcal{L}_{\text{total}}$
 4.2 Design Details
@@
 where $\boldsymbol{x}_{a}^{j,i}$ denotes the adversarial example generated by the $i$-th base attack at stage $j$.
-The softmax operation ensures that the final combination remains within the original $\ell_{p}$-norm constraints, as we aim to preserve the structure of the base attacks without modifying them directly.
-This linear formulation provides a simple and extensible mechanism for combining base attacks.
+This linear formulation, with softmax probabilities providing normalized importance weights, provides a simple and extensible mechanism for combining base attacks while preserving the overall perceptual similarities.
 Future work may explore more advanced combination strategies beyond weighted averaging, or treat individual hyperparameters of base attacks, which are fixed in this work, as additional learnable variables.
@@
 This enables iterative refinement of the adversarial perturbation over multiple stages.
-While we adopt a simple linear chaining strategy here, more sophisticated chaining mechanisms (e.g., cyclic or attention) are possible and left as future work.
-Finally, we define the meta-loss that guides the optimization of the attack weights: $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{ssim}}\cdot(1-\text{SSIM}(\boldsymbol{x},\boldsymbol{x}_{a}))$ The first term encourages attack effectiveness by maximizing the confidence drop in the ground-truth class $y$ for the adversarial example $\boldsymbol{x}_{a}$.
+While we adopt a linear chaining strategy here, other chaining mechanisms (e.g., cyclic or attention) are possible and left as future work.
+Finally, we define the meta-loss that guides the optimization of the attack weights: $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{ssim}}\cdot(1-\text{SSIM}(\boldsymbol{x},\boldsymbol{x}_{a}))$.
+The first term encourages attack effectiveness by maximizing the confidence drop in the ground-truth class $y$ for the adversarial example $\boldsymbol{x}_{a}$.
 The second term penalizes perceptual distortion, measured using the Structural Similarity Index Measure (SSIM), which ranges from 0 to 1, with higher values indicating greater similarity.
+We include SSIM in the meta-loss to constraint DASH adversarial examples within a visual perturbation budget.
+While it is not possible to strictly clip an example to a specific SSIM/FID/LPIPS budget limit, the computation of these metrics remains differentiable and hence can be used as an optimization parameter.
+In the experiment, the weights are optimized until some threshold SSIM is not achieved.
 In practice, we observe that substituting SSIM with other perceptual similarity metrics yields comparable results.
 Our full attack algorithm is summarized in Algorithm 1 and a visual demonstration is also shown in Figure 2.
-Algorithm 1 Search for Optimal Attack Sequence $\mathcal{A}_{\Theta}$
-1:Randomly Initialize $\boldsymbol{\alpha}$
-[…]
-3:for $\text{epoch}=1$ to $T$ do
-4:  Sample a mini-batch $\{(\boldsymbol{x},y)\}\subset\mathcal{D}$
-5:  Generate adversarial examples: $\boldsymbol{x}_{a}\leftarrow\mathcal{A}(\boldsymbol{x},y)$
-6:  Feed to target model: $f(\boldsymbol{x}_{a})$
-7:  Compute total loss $\mathcal{L}_{\text{total}}$
-8:  Update parameters: $\boldsymbol{\alpha}\leftarrow\boldsymbol{\alpha}-\eta\,\nabla_{\boldsymbol{\alpha}}\mathcal{L}_{\text{total}}(\boldsymbol{x}_{a},y)$
-9:end for
-10:Return attack sequence $\mathcal{A}_{\Theta}$ that minimizes $\mathcal{L}_{\text{total}}$
-Table 1: Performance comparison of adversarial attacks across different datasets and RobustBench models.
-Base is the ASR without applying any post-processing defenses; Avg.
+5 Experiments and Results
+Next we discuss our core experimental results.
+All results are obtained using NVIDIA A100 GPU (unless specified).
+Significant amount of additional experimental results and details (e.g.
+black-box attacks as DASH base attacks, DASH attack weight distributions, base attack pool variations) can be found inside the supplementary document (appendix).
+Source code for DASH is also provided as supplementary materials.
+5.1 Experimental Setup
+We discuss the datasets, robust models, post-processing defenses, metrics, base attacks, and DASH training.
+5.1.1 Datasets and Robust Models
+Our experiments are conducted on three commonly used benchmark datasets of different scales: CIFAR-10 [krizhevsky2014cifar], CIFAR-100 [Krizhevsky09learningmultiple], and ImageNet [deng2009imagenet].
+For each dataset, we randomly sample 1,000 test images.
+Since standard models without explicit defenses are easily broken by existing attacks, our evaluation focuses on robust models curated by RobustBench [croce2020robustbench], a widely used benchmark for adversarial robustness in which the models are variants of standard adversarial training [madry2017towards].
+Table 1: Performance of attacks across datasets and RobustBench [croce2020robustbench] models.
+Base is the ASR without any post-processing defenses; Avg.
 is the average of all the ASR values.
 SSIM values are scaled to [0,100].
-Dataset Target Model Attack ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
+Time (sec) is the average generation time of one example.
+Dataset Target Model Attack Time (s) ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
 Base JPEG TVM Ensemble NRP Avg.
-CIFAR-100 Cui et al. (2024) DI-FGSM 65.20 65.49 65.35 65.55 57.03 63.72 92.08 0.0168 54.93
-CW 100.00 87.70 96.68 78.61 57.71 84.14 94.38 0.0153 32.49
-PI-FGSM++ 64.41 65.78 64.32 65.64 50.86 62.20 91.12 0.0199 58.94
-AutoAttack 82.56 81.33 79.51 80.55 59.88 76.77 92.09 0.0167 41.68
-DiffPGD 76.20 76.00 75.80 75.60 75.40 75.80 91.15 0.0165 43.50
-DiffAttack 77.50 77.20 77.70 77.40 76.40 77.24 91.16 0.0169 42.30
-AdvAD 82.20 82.80 81.10 81.70 67.90 79.14 83.18 0.0288 45.67
+CIFAR-100 Cui2024 [cui2024decoupled] DI-FGSM 0.17 65.20 65.49 65.35 65.55 57.03 63.72 92.08 0.0168 54.93
+CW 0.11 100.00 87.70 96.68 78.61 57.71 84.14 94.38 0.0153 32.49
+PI-FGSM++ 0.22 64.41 65.78 64.32 65.64 50.86 62.20 91.12 0.0199 58.94
+AutoAttack 0.33 82.56 81.33 79.51 80.55 59.88 76.77 92.09 0.0167 41.68
+DiffPGD 4.21 76.20 76.00 75.80 75.60 75.40 75.80 91.15 0.0165 43.50
+DiffAttack 12.44 77.50 77.20 77.70 77.40 76.40 77.24 91.16 0.0169 42.30
+AdvAD 0.51 82.20 82.80 81.10 81.70 67.90 79.14 83.18 0.0288 45.67
 […]
-Wang et al. (2023) DI-FGSM 66.46 67.01 66.07 66.72 59.86 65.22 91.96 0.0169 55.24
-CW 100.00 85.74 95.31 76.95 59.86 83.57 94.55 0.0161 32.19
-PI-FGSM++ 66.17 67.21 65.29 66.76 52.30 63.55 91.14 0.0199 57.88
-AutoAttack 83.50 82.91 81.05 81.64 62.60 78.34 91.25 0.0189 42.19
-DiffPGD 79.10 76.10 74.20 73.80 73.40 75.32 91.50 0.0161 51.80
-DiffAttack 76.90 76.60 76.40 76.80 76.30 76.53 90.90 0.0176 42.20
-AdvAD 79.90 79.60 79.90 79.50 67.20 77.42 83.07 0.0293 47.29
+Wang2023 [wang2023better] DI-FGSM 0.17 66.46 67.01 66.07 66.72 59.86 65.22 91.96 0.0169 55.24
+CW 0.12 100.00 85.74 95.31 76.95 59.86 83.57 94.55 0.0161 32.19
+PI-FGSM++ 0.22 66.17 67.21 65.29 66.76 52.30 63.55 91.14 0.0199 57.88
+AutoAttack 0.39 83.50 82.91 81.05 81.64 62.60 78.34 91.25 0.0189 42.19
+DiffPGD 4.77 79.10 76.10 74.20 73.80 73.40 75.32 91.50 0.0161 51.80
+DiffAttack 12.96 76.90 76.60 76.40 76.80 76.30 76.53 90.90 0.0176 42.20
+AdvAD 0.51 79.90 79.60 79.90 79.50 67.20 77.42 83.07 0.0293 47.29
 […]
-(Addepalli et al.) (2022) DI-FGSM 72.09 72.13 71.84 71.74 63.42 70.24 92.05 0.0156 56.69
-CW 100.00 87.21 96.48 77.93 60.74 84.47 95.56 0.0113 26.94
-PI-FGSM++ 71.50 71.93 70.86 71.15 56.89 68.47 91.14 0.0192 59.94
-AutoAttack 83.77 82.70 81.02 80.68 61.33 77.90 93.03 0.0147 39.59
-DiffPGD 82.50 82.20 81.80 81.60 80.90 81.80 91.05 0.0168 42.80
-DiffAttack 81.60 81.60 81.40 81.40 81.50 81.35 91.26 0.0163 42.15
-AdvAD 84.50 84.00 83.70 84.30 70.50 81.60 83.92 0.0291 45.63
+Addepalli2022 [addepalli2022efficient] DI-FGSM 0.18 72.09 72.13 71.84 71.74 63.42 70.24 92.05 0.0156 56.69
+CW 0.13 100.00 86.67 96.45 79.04 58.72 84.18 94.94 0.0118 27.54
+PI-FGSM++ 0.23 71.50 71.93 70.86 71.15 56.89 68.47 91.14 0.0192 59.94
+AutoAttack 0.39 83.77 82.70 81.02 80.68 61.33 77.90 93.03 0.0147 39.59
+DiffPGD 4.27 82.50 82.20 81.80 81.60 80.90 81.80 91.05 0.0168 42.80
+DiffAttack 12.86 81.60 81.60 81.40 81.40 81.50 81.35 91.26 0.0163 42.15
+AdvAD 0.54 84.50 84.00 83.70 84.30 70.50 81.60 83.92 0.0291 45.63
 […]
-CIFAR-10 (Bartoldson et al.) (2024) DI-FGSM 41.39 42.46 42.27 42.75 35.70 40.91 90.88 0.0200 59.07
-CW 97.27 61.62 96.09 60.35 58.59 74.78 95.25 0.0206 55.25
-PI-FGSM++ 45.80 48.46 46.72 48.89 23.24 42.62 88.74 0.0278 66.91
-AutoAttack 73.24 71.88 72.75 71.19 39.36 65.68 87.51 0.0264 53.05
-DiffPGD 53.20 52.40 52.60 52.40 52.80 52.68 91.60 0.0165 54.20
-DiffAttack 73.60 73.30 73.60 73.50 72.90 73.38 82.90 0.0369 51.60
-AdvAD 61.30 60.70 60.60 61.20 41.10 56.98 83.50 0.0269 49.37
+CIFAR-10 Bartold2024 [bartoldson2024adversarial] DI-FGSM 0.72 41.39 42.46 42.27 42.75 35.70 40.91 90.88 0.0200 59.07
+CW 0.62 97.27 61.62 96.09 60.35 58.59 74.78 95.25 0.0206 55.25
+PI-FGSM++ 0.83 45.80 48.46 46.72 48.89 23.24 42.62 88.74 0.0278 66.91
+AutoAttack 2.17 73.24 71.88 72.75 71.19 39.36 65.68 87.51 0.0264 53.05
+DiffPGD 4.80 53.20 52.40 52.60 52.40 52.80 52.68 91.60 0.0165 54.20
+DiffAttack 15.91 73.60 73.30 73.60 73.50 72.90 73.38 82.90 0.0369 51.60
+AdvAD 3.99 61.30 60.70 60.60 61.20 41.10 56.98 83.50 0.0269 49.37
 […]
-Cui et al. (2024) DI-FGSM 51.64 52.71 52.36 53.01 45.23 50.99 90.51 0.0206 62.20
-CW 99.80 67.68 95.51 65.33 65.33 78.73 95.39 0.0206 51.38
-PI-FGSM++ 55.76 57.09 56.05 56.84 31.07 51.36 88.71 0.0272 68.81
-AutoAttack 78.42 77.34 78.22 76.46 44.34 70.96 86.88 0.0273 56.71
-DiffPGD 76.20 75.80 76.50 76.10 75.40 76.00 79.80 0.0350 60.50
-DiffAttack 82.80 82.50 83.10 82.70 82.20 82.66 81.36 0.0404 58.24
-AdvAD 66.80 67.20 66.60 66.20 45.50 62.46 83.11 0.0265 50.95
+Cui2024 [cui2024decoupled] DI-FGSM 0.16 51.64 52.71 52.36 53.01 45.23 50.99 90.51 0.0206 62.20
+CW 0.11 99.80 67.68 95.51 65.33 65.33 78.73 95.39 0.0206 51.38
+PI-FGSM++ 0.21 55.76 57.09 56.05 56.84 31.07 51.36 88.71 0.0272 68.81
+AutoAttack 0.41 78.42 77.34 78.22 76.46 44.34 70.96 86.88 0.0273 56.71
+DiffPGD 5.30 76.20 75.80 76.50 76.10 75.40 76.00 79.80 0.0350 60.50
+DiffAttack 12.88 82.80 82.50 83.10 82.70 82.20 82.66 81.36 0.0404 58.24
+AdvAD 0.54 66.80 67.20 66.60 66.20 45.50 62.46 83.11 0.0265 50.95
 […]
-Wang et al. (2023) DI-FGSM 51.64 52.13 52.09 51.84 44.57 50.45 90.45 0.0203 63.21
-CW 98.90 70.21 94.82 66.11 51.76 76.36 95.25 0.0208 52.86
-PI-FGSM++ 55.29 57.11 54.61 56.27 29.61 50.58 88.72 0.0265 70.71
-AutoAttack 77.93 77.54 77.73 76.27 43.16 70.53 86.83 0.0276 56.75
-DiffPGD 75.80 75.40 76.10 75.60 75.10 75.60 79.50 0.0345 60.80
-DiffAttack 82.50 82.40 82.50 82.40 82.10 82.38 80.12 0.0429 57.22
-AdvAD 67.40 68.40 66.40 67.20 44.10 62.70 82.64 0.0271 52.78
+Wang2023 [wang2023better] DI-FGSM 0.17 51.64 52.13 52.09 51.84 44.57 50.45 90.45 0.0203 63.21
+CW 0.12 98.90 70.21 94.82 66.11 51.76 76.36 95.25 0.0208 52.86
+PI-FGSM++ 0.22 55.29 57.11 54.61 56.27 29.61 50.58 88.72 0.0265 70.71
+AutoAttack 0.41 77.93 77.54 77.73 76.27 43.16 70.53 86.83 0.0276 56.75
+DiffPGD 5.13 75.80 75.40 76.10 75.60 75.10 75.60 79.50 0.0345 60.80
+DiffAttack 13.49 82.50 82.40 82.50 82.40 82.10 82.38 80.12 0.0429 57.22
+AdvAD 0.54 67.40 68.40 66.40 67.20 44.10 62.70 82.64 0.0271 52.78
 […]
-ImageNet-1k Salman et al. (2020) DI-FGSM 83.90 83.60 83.00 82.60 75.40 81.70 91.42 0.0849 49.39
-CW 98.50 88.40 98.20 76.00 63.70 85.02 97.75 0.0263 6.77
-PI-FGSM++ 87.10 86.80 85.60 85.10 74.50 83.82 90.54 0.0960 49.57
-AutoAttack 94.90 94.30 91.30 90.30 72.60 91.56 92.90 0.0757 21.78
-DiffPGD 76.20 69.40 74.80 68.50 72.30 72.24 87.50 0.0892 45.60
-DiffAttack 92.50 86.80 88.90 87.90 89.80 89.18 89.17 0.0763 40.78
-AdvAD 89.20 87.90 87.70 87.90 86.80 87.90 71.30 0.1949 50.14
+ImageNet-1k Salman2020 [salman2020adversarially] DI-FGSM 1.22 83.90 83.60 83.00 82.60 75.40 81.70 91.42 0.0849 49.39
+CW 0.58 98.50 88.40 98.20 76.00 63.70 85.02 97.75 0.0263 6.77
+PI-FGSM++ 1.75 87.10 86.80 85.60 85.10 74.50 83.82 90.54 0.0960 49.57
+AutoAttack 2.22 94.90 94.30 91.30 90.30 72.60 91.56 92.90 0.0757 21.78
+DiffPGD 10.40 76.20 69.40 74.80 68.50 72.30 72.24 87.50 0.0892 45.60
+DiffAttack 20.33 92.50 86.80 88.90 87.90 89.80 89.18 89.17 0.0763 40.78
+AdvAD 5.61 89.20 87.90 87.70 87.90 86.80 87.90 71.30 0.1949 50.14
 […]
-5 Experiments and Results
-Next we discuss our experimental setup and evaluations. All the reported results are obtained using NVIDIA A100 GPU.
-5.1 Experimental Setup
-5.1.1 Datasets and Models
-Our experiments are conducted on three commonly used benchmark datasets of different scales: CIFAR-10 Krizhevsky et al.
-(2014), CIFAR-100 Krizhevsky (2009), and ImageNet Deng et al.
-(2009).
-For each dataset, we randomly sample 1,000 test images that are correctly classified by the victim model.
-Since standard models (without any defenses) are easily broken by existing attacks, we focus our evaluation on robust models curated by RobustBench Croce et al.
-(2020), a popular benchmark for adversarial robustness.
 5.1.2 Post-Processing Defenses
-To further test the limits of DAASH and other SOTA attacks in terms of generating potent adversarial examples, we also experiment with post-processing defenses (on top of using RobustBench models).
-The post-processing defenses include JPEG compression Guo et al.
-(2017), Total Variation Minimization (TVM) Wang et al.
-(2018), Neural Representation Purifier (NRP) Naseer et al.
-(2020), and an ensemble of four methods: bit-depth reduction Xu (2018), JPEG, TVM, and Non-Local Means Denoising (NLM) Buades et al.
-(2011).
+To further test the limits of DASH and other SOTA attacks in terms of generating potent adversarial examples, we also experiment with post-processing defenses (on top of using RobustBench models).
+The post-processing defenses include JPEG compression [guo2017countering], Total Variation Minimization (TVM) [wang2018adversarial], Neural Representation Purifier (NRP) [naseer2020self], and an ensemble of four methods: bit-depth reduction [xu2018feature], JPEG, TVM, and Non-Local Means Denoising (NLM) [buades2011non].
+Table 2: Transferability Comparison on CIFAR-100 Dataset. S represents Surrogate models and T represents Target models.
+S (↓) T (→) Attacks [cui2024decoupled] [wang2023better] [addepalli2022efficient] Avg.(↑) SSIM (↑)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+CW 37.32 $73 \cellcolorgray$!1584.18 53.08 94.94
+[…]
+[…]
+[…]
+[…]
 5.1.3 Evaluation Metrics
 We use attack success rate (ASR) to measure attack potency, and for determining human visual imperceptibility, we use: SSIM, LPIPS, and FID.
-Figure 3: Comparison of DAASH performance before and after training, illustrating ASR and SSIM across different defense settings.
-“Trained DAASH" significantly outperforms the randomly weighted version (“Random DAASH").
-5.1.4 Base Attacks and DAASH Training Details
-We include 10 diverse base attacks in our attack pool: FGSM Goodfellow et al.
-(2015b), PGD Madry et al.
-(2018b), CW Carlini and Wagner (2017), FAB Croce and Hein (2020a), TI-FGSM Dong (2019), NI-FGSM Lin et al.
-(2019), MI-FGSM Dong et al.
-(2017), DI-FGSM Xie et al.
-(2018), BIM Kurakin et al.
-(2016), and PI-FGSM++ Gao et al.
-(2020).
+Figure 3: Comparison of DASH performance before and after training, illustrating ASR and SSIM across different defense settings.
+“Trained DASH” significantly outperforms the randomly weighted version (“Random DASH”).
+5.1.4 Base Attacks and DASH Training Details
+We include 10 diverse base attacks in our attack pool: FGSM [goodfellow2015explaining], PGD [madry2018towards], CW [carlini2017towards], FAB [croce2020minimally], TI-FGSM [dong2019evading], NI-FGSM [lin2019nesterov], MI-FGSM [dong2017boosting], DI-FGSM [xie2018improving], BIM [kurakin2016adversarial], and PI-FGSM++ [gao2020patch].
 Each attack offers unique advantages and covers a broad range of perturbation strategies.
-In addition, we include a None operation at each stage, which simply returns the input image, allowing the model to flexibly decide whether to skip processing in a given stage.
-For the meta-loss, we set $\lambda_{\text{asr}}=1.3$ and $\lambda_{\text{ssim}}=1.0$, which work well across datasets.
-We run Algorithm 1 for $T=100$ epochs using the Adam optimizer with a learning rate of 0.001.
-After training, the learned weights for combining base attacks in each stage are used to generate the final adversarial example.
-Figure 4: DAASH training progression over epochs.
-5.2 Effectiveness of DAASH against Robust Models
-Table 1 reports the attack success rate (ASR) and perceptual similarity scores of DAASH and seven representative state-of-the-art attacks evaluated against seven recent adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet.
-As shown in the table, DAASH consistently outperforms all baseline attacks in terms of ASR under both the Base (no post-processing) and post-processing defense conditions, across all datasets.
-This includes outperforming recent attacks explicitly designed to go beyond $\ell_{p}$-norm constraints.
-For example, on CIFAR-100, the best-performing baseline is CW with an average ASR of $84\%$, while the most recent perceptual-aligned attack, AdvAD, achieves only $79\%$.
-In contrast, DAASH achieves an ASR close to $100\%$.
-Moreover, DAASH yields better imperceptibility scores compared to these dedicated perceptual attacks.
-Similar trends are observed on CIFAR-10 and ImageNet: DAASH improves over the best baseline by at least $15\%$ ASR on CIFAR-10 and $5\%$ on ImageNet, on average.
-Its perceptual scores are also consistently better than other attacks specifically designed for perceptual alignment.
-Table 2: Effectiveness of DAASH with varying number of stages for the robust CIFAR-100 model – Cui et al. (2024).
+For $\ell_{\infty}$-norm attacks, we set the perturbation budget to $\epsilon=\frac{7}{255}$.
+For different variants of FGSM evaluated in this paper, we use a momentum factor of 1.
+In TI-FGSM [dong2019evading] and DI-FGSM [xie2018improving], we set the resize_rate to 0.9, diversity_prob to 0.7, and random_start to False.
+In PI-FGSM++ [gao2020patch], we set the probability of using diverse inputs to 0.7 and the project_factor to 0.8.
+For the Fast Adaptive Boundary (FAB) attack [croce2020minimally], we set $\alpha_{\text{max}}=0.1$, overshooting factor $\eta=1.05$, and the backward step parameter $\beta=0.9$.
+For BIM [kurakin2016adversarial] and PGD [madry2017towards], we use a step size of $\alpha=0.01$ and number of iterations $T=10$.
+For the CW attack [carlini2017towards], we set the confidence parameter $c=10$, margin parameter $\kappa=0$, and optimized for 100 steps using a learning rate of $0.01$.
+In addition, we include a None operation at each stage, which returns the input image, allowing the model to flexibly decide whether to skip processing in a given stage.
+Table 3: Effectiveness of DASH with varying numbers of stages for the robust CIFAR-100 Cui2024 [cui2024decoupled] model.
 Num. of ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
@@
 5 100.00 100.00 100.00 100.00 99.90 99.98 92.92 0.0209 48.25
-Table 3: Average attack success rate (ASR↑) and structural similarity (SSIM↑) for cross-model transferability.
+For the meta-loss, we set $\lambda_{\text{asr}}=1.3$ and $\lambda_{\text{ssim}}=1.0$, which work well across datasets.
+We run Algorithm 1 for $T=100$ epochs using the Adam [kingma2014adam] optimizer with a learning rate of 0.01.
+After training, the learned weights for combining base attacks in each stage are used to generate the final adversarial example.
+A visualization of this attack combination has been added in the supplimentary material.
+Figure 4: DASH training progression over epochs.
+5.2 Effectiveness of DASH against Robust Models
+Table 4: Average attack success rate (ASR↑) and structural similarity (SSIM↑) for cross-model weight transferability.
 Trained Tested CIFAR-100 CIFAR-10 ImageNet-1K
-(Cui et al.) (2024) (Wang et al.) (2023) (Addepalli et al.) (2022) (Bartoldson et al.) (2024) (Cui et al.) (2024) (Wang et al.) (2023) (Salman et al.) (2020)
+Cui2024 [cui2024decoupled] Wang2023 [wang2023better] Addepalli22 [addepalli2022efficient] Bartold2024 [bartoldson2024adversarial] Cui2024 [cui2024decoupled] Wang2023 [wang2023better] Salman20 [salman2020adversarially]
 Dataset ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM ASR SSIM
@@
 […]
+Table 1 reports the attack success rate (ASR) and perceptual similarity scores of DASH and seven representative state-of-the-art attacks evaluated against seven recent adversarially trained robust models across CIFAR-10, CIFAR-100, and ImageNet.
+As shown in Table 1, DASH consistently outperforms all baseline attacks in terms of ASR under both the Base (no post-processing) and post-processing defense conditions, across all datasets.
+This includes outperforming recent attacks explicitly designed to go beyond $\ell_{p}$-norm constraints.
+For example, on CIFAR-100, the best-performing baseline is CW with an average ASR of $84\%$, while the most recent perceptual-aligned attack, AdvAD, achieves only $79\%$.
+In contrast, DASH achieves an ASR close to $100\%$.
+Moreover, DASH yields better imperceptibility scores compared to these dedicated perceptual attacks.
+Similar trends are observed on CIFAR-10 and ImageNet: DASH improves over the best baseline by at least $15\%$ ASR on CIFAR-10 and $5\%$ on ImageNet, on average.
+Its perceptual scores are also consistently better than other attacks specifically designed for perceptual alignment.
+We have obtained results with different parameter configurations for each compared attacks and reported the best results.
 Effectiveness of CW.
@@
 This suggests that some classical attacks, when properly tuned, can potentially still be highly effective in terms of both ASR and imperceptibility.
-Although DAASH outperforms CW in terms of ASR, its imperceptibility scores are sometimes slightly lower.
-Intuitively, since CW is one of the base attacks in DAASH, it could be expected to receive higher attention weights during optimization.
+Although DASH outperforms CW in terms of ASR, its imperceptibility scores are sometimes slightly lower.
+Intuitively, since CW is one of the base attacks in DASH, it could be expected to receive higher attention weights during optimization.
 However, this is not always the case.
 The reason lies in the mismatch between the objectives: CW uses a hinge-like loss that stops contributing gradients once misclassification is achieved, whereas our meta-loss continues to penalize the confidence of the true class beyond the decision boundary.
-As a result, after initial misclassification, CW receives lower weights in DAASH, which shifts focus to other attacks that still produce informative gradients.
-This continued “pushing” of adversarial examples beyond the boundary enhances robustness of DAASH against additional defenses such as post-processing but comes at the cost of slightly reduced perceptual similarity.
-5.3 Ablation Studies on DAASH
-Our DAASH framework consists of two key components: (1) learnable weights for each base attack within every attack cell, and (2) chained attack cells to better escape poor local optima.
+As a result, after initial misclassification, CW receives lower weights in DASH, which shifts focus to other attacks that still produce informative gradients.
+This continued “pushing” of adversarial examples beyond the boundary enhances robustness of DASH against additional defenses such as post-processing where CW fails.
+5.3 Effectiveness of DASH on Black-Box Settings
+Table 2 reports the attack performances of DASH in black-box scenarios.
+Additional preliminary results of using black-box attacks (e.g.
+Sign Flip [chen2020boosting], Rays [chen2020rays], Square [andriushchenko2020square], Sign Hunter [al2020sign]) as DASH base attacks can be found in the supplementary document (appendix).
+Adversarial Examples from Surrogate Model.
+Table 2 shows that the adversarial examples generated with DASH for one model also perform well for other models.
+DASH achieves the strongest transferability, achieving the top average ASR for each model while also maintaining the best perceptual quality (SSIM).
+For Cui2024 [cui2024decoupled] and Wang2023 [wang2023better] models DASH outperforms the best baseline Diffattack with 14% and 16% improvement in average ASR, keeping the best SSIM score.
+For Addepalli2022 [addepalli2022efficient] DASH outperforms the best baseline AdvAD with almost a 5% improvement in average ASR, keeping the best SSIM score.
+This indicates that DASH consistently achieves good result in both transferability and imperceptibility scenarios.
+5.4 Ablation Studies on DASH
+Our DASH framework consists of two key components: (1) learnable weights for each base attack within every attack cell, and (2) chained attack cells to better escape poor local optima.
 In this section, we conduct ablation studies to demonstrate the importance of each component.
 Importance of Learnable Weights.
-To evaluate the benefit of using learnable weights, we compare the performance of DAASH against a variant where the weights for base attacks are initialized randomly and left untrained.
+To evaluate the benefit of using learnable weights, we compare the performance of DASH against a variant where the weights for base attacks are initialized randomly and left untrained.
 Specifically, the random weights are sampled uniformly in the range $[0,1]$ and used directly without optimization.
 We repeat this random setup 100 times and report the average results.
-Figure 3 shows that using trained weights significantly improves both ASR and imperceptibility (measured by SSIM) across all types of robust models.
+Figure 3 shows that using trained weights significantly improves both ASR and imperceptibility (SSIM) across all types of robust models.
 Furthermore, Figure 4 illustrates the ASR and SSIM trends as training progresses over epochs.
-Both metrics exhibit clear improvement—ASR increases across all settings, and SSIM improves almost monotonically—highlighting the effectiveness of our meta-learning process.
-These results confirm that the learnable weighting mechanism in DAASH plays a crucial role.
+Both metrics exhibit clear improvement: ASR increases across all settings, and SSIM improves almost monotonically, highlighting the effectiveness of our meta-learning process.
+These results confirm that the learnable weighting mechanism in DASH plays a crucial role.
 Importance of Multiple Stages.
-To assess the value of chaining multiple stages, we evaluate DAASH with varying numbers of stages.
+To assess the value of chaining multiple stages, we evaluate DASH with varying numbers of stages.
 Table 3 reports the ASR for different stage configurations.
-The results show a substantial ASR jump when increasing from one to two stages (from 51.88% to 99.49%), suggesting that two stages are sufficient to escape bad local optima and refine the adversarial example generation.
-Adding more stages yields marginal improvements, indicating diminishing returns beyond two stages.
+The results show a substantial ASR jump when increasing from one to two stages (from 51.88% to 99.49%), suggesting that multi-stage can help escape bad local optima and refine the adversarial example generation.
+Adding more stages may yield improvements.
 Effectiveness against Unseen Defenses.
-We also examine whether the learned weights in DAASH are transferable across models and datasets.
+We also examine whether the learned weights in DASH are transferable across models and datasets.
 Table 4 shows that weights trained on one model or dataset can be applied to unseen models or datasets while still achieving near-optimal results.
 The variation in average ASR remains within 3%, and SSIM varies by less than 2.
-These results demonstrate that DAASH can generalize well across different settings/models.
+These results demonstrate that DASH can generalize well across different settings/models.
 6 Conclusion
-In this work, we introduced DAASH, a novel differentiable framework for learning optimal compositions of adversarial attacks in a white-box setting.
-Unlike traditional handcrafted methods that heuristically combine attacks, DAASH leverages gradient-based optimization to learn soft weightings over multiple attack modules across multiple stages.
-This process is guided by a meta-loss that jointly captures attack effectiveness and perceptual alignment with human vision.
-Our flexible framework enables the transformation of existing norm-constrained attacks into stronger attacks that also produce perceptually aligned adversarial examples.
-Empirical results demonstrate that DAASH  significantly outperforms state-of-the-art perceptually aligned attacks in both attack success rate and imperceptibility scores.
-DAASH  generalizes well to unseen defenses, establishing itself as a strong and adaptive baseline for future robustness evaluation.
+In this work, we introduced DASH, a novel differentiable framework for learning optimal compositions of adversarial attacks in both white-box and black-box settings.
+Unlike traditional methods that heuristically combine attacks, DASH leverages gradient-based optimization to learn soft weightings over multiple attack modules across stages.
+DASH is guided by a meta-loss that jointly captures attack effectiveness and perceptual alignment with human vision.
+Our flexible framework enables the transformation of existing norm-constrained attacks into stronger attacks that also produce perceptually aligned and transferable adversarial examples.
+Empirical results demonstrate that DASH  significantly outperforms state-of-the-art perceptually aligned attacks in attack success rate, transferability and imperceptibility scores.
+DASH  generalizes well to unseen defenses, establishing itself as a strong and adaptive baseline for future robustness evaluation.
 Overall, our results underscore the potential of differentiable search in the adversarial example literature and open up new directions for developing more adaptive and powerful attack strategies.
-Future work may explore extending DAASH  to black-box or query-limited scenarios, as well as integrating it with generative models to further enhance imperceptibility and increase sample diversity.
+Future work may explore extending DASH for attacking multimodal models and generative tasks.
+Appendix A Appendix
+Below, we provide additional details on the base attacks, defense strategies, and evaluation metrics used in our study (Appendix B).
+We then present the hyperparameters for each base attack and detail how they are selected across different stages of DASH (Appendix C).
+The following sections demonstrate the extension of the DASH framework to black-box settings (Appendix D) and examine the performance of DASH when optimizing with various base attacks and imperceptibility metrics (Appendix E).
+Table 5: Performance of DASH with Black-Box attacks for CIFAR-10 Dataset.
+ASR ( $\uparrow$ ) SSIM (↑) LPIPS (↓) FID (↓)
+Model Attack Method Base JPEG TVM Ensemble NRP Avg.
+ResNet-18 [he2016deep] Sign Flip [chen2020boosting] 93.07 25.68 73.44 26.86 27.34 49.28 93.50 0.0025 58.06
+Rays [chen2020rays] 94.53 55.08 48.54 43.26 28.03 53.89 94.31 0.0120 32.37
+Square [andriushchenko2020square] 87.89 56.25 79.98 43.85 29.59 59.51 94.82 0.0160 34.87
+Sign Hunter [al2020sign] 84.18 57.91 78.81 50.88 27.93 59.94 94.32 0.0174 35.54
+[…]
+ViT [dosovitskiy2020image] Sign Flip [chen2020boosting] 93.65 36.04 73.83 30.96 31.54 53.20 92.09 0.0037 64.83
+Rays [chen2020rays] 91.70 63.18 61.13 54.79 38.48 61.86 94.52 0.0293 33.81
+Square [andriushchenko2020square] 92.29 62.40 84.18 47.36 46.29 66.50 96.22 0.0106 31.71
+Sign Hunter [al2020sign] 93.36 69.24 87.11 57.62 38.87 69.24 95.79 0.0177 30.19
+[…]
+Appendix B Additional Background
+Figure 5: Learned DASH weights (represented as % of contribution) for different base attacks across stages.
+B.1 Norm-Constrained Attacks
+Classical adversarial methods constrain perturbations using $\ell_{p}$ norms by either enforcing them directly [madry2017towards] or indirectly [carlini2017towards].
+PGD Attack The PGD attack seeks an adversarial example $\boldsymbol{x}_{a}$ such that $\|\boldsymbol{x}_{a}-\boldsymbol{x}\|_{\infty}\leq\epsilon$, where $\epsilon$ denotes the hard norm constraint and we denote the norm-constraint as $B_{\epsilon}$.
+and $\boldsymbol{x}_{a}$ maximizes a loss $\mathcal{L}$ such as the cross-entropy loss.
+The update process is iterative.
+The first iteration is initialized as $\boldsymbol{x}_{a}^{(0)}=x+\delta_{0},~~\delta_{0}\sim\mathcal{U}(-\epsilon,\epsilon)$ and is later updated in step $t$ as
+$\boldsymbol{x}_{a}=\Pi_{B_{\epsilon}}\left(\boldsymbol{x}_{a}^{(t)}+\alpha\cdot\text{sign}\left(\nabla_{x}\mathcal{L}(\boldsymbol{x}_{a}^{(t)},y;\theta)\right)\right),$
+where $\alpha$ denotes the step size in each iteration, $\Pi$ denotes the projection operation onto the ball $B_{\epsilon}$, centered at $\boldsymbol{x}_{a}$.
+The attack runs a total of $T$ iterations.
+FGSM is a single step update:
+$\boldsymbol{x}_{a}=\Pi_{B_{\epsilon}}\left(\boldsymbol{x}+\epsilon\cdot\text{sign}\left(\nabla_{x}\mathcal{L}(\boldsymbol{x}_{a}^{(t)},y;\theta)\right)\right).$
+Both the FGSM and PGD attack can be further enhanced, mostly for the purpose of improved transferability to unknown models without white-box access, by provideing enhanced inputs to the the optimization process [xie2018improving, gao2020patch].
+CW Attack The Carlini & Wagner (CW) [carlini2017towards] attack enforces the norm constraints softly and formulates adversarial example generation as the following optimization without explicit projections:
+$\min_{\boldsymbol{\delta}}\|\boldsymbol{\delta}\|_{2}^{2}+c\cdot\max(0,Z(\boldsymbol{x}+\delta)_{y}-\max_{i\neq y}Z(\boldsymbol{x}+\boldsymbol{\delta})_{i}+\kappa)$ (4)
+where $Z(x)$ represents pre-softmax logits, $c$ balances the two objectives, and $\kappa$ controls confidence margins, which is usually set 0 for common adversarial example generation.
+Table 6: Performance comparison of DASH with varying base attacks for CIFAR-100 Cui2024 [cui2024decoupled] model.
+Base Attacks ASR ( $\uparrow$ ) SSIM ( $\uparrow$ ) LPIPS ( $\downarrow$ ) FID ( $\downarrow$ ) W/O Def.
+JPEG TVM Ensemble NRP Avg PGD, MI-FGSM, CW 99.90 97.75 99.90 95.90 97.36 98.16 90.75 0.0191 60.31 NI-FGSM, CW, FGSM, DI-FGSM, None 99.41 94.82 99.32 93.16 94.73 96.29 93.33 0.0146 49.61 TI-FGSM, PGD, PI-FGSM++, CW, MI-FGSM, FGSM, None 100.00 99.90 100.00 99.51 99.71 99.82 93.90 0.0133 45.73 CW, BIM, NI-FGSM, PGD, DI-FGSM, TI-FGSM, PI-FGSM++, FGSM, MI-FGSM 97.75 94.34 97.46 92.77 94.82 95.43 92.35 0.0193 51.24 None, FGSM, PGD, CW, FAB, TI-FGSM, NI-FGSM, MI-FGSM, DI-FGSM, BIM, PI-FGSM++ 100.00 99.90 100.00 99.32 99.61 99.77 94.43 0.0139 41.08
+B.2 Perceptually-Informed Attack Design
+Moving beyond norm-constrained attacks, recent methods incorporate perceptual metrics directly into optimization objectives to improve their alignments with human vision.
+B.2.1 Frequency-Domain Approaches
+Recognizing human visual system characteristics, frequency-based methods constrain perturbations to less perceptible components.
+Semantic Similarity Attack on High-frequency (SSAH) [luo2022frequency] utilizes discrete wavelet transforms to penalize low-frequency modifications.
+AdvDrop [duan2021advdrop] applies perturbations in DCT space, dropping high-frequency components that humans cannot perceive.
+B.2.2 Perceptual Distance Integration
+PerC-AL [zhao2020towards] incorporates perceptual color distance to measure the perturbation magnitude and adds this metric as an additional term in the optimization process:
+$\mathcal{L}_{per}=\sum_{i,j}|\Delta E_{ij}(x,x_{adv})|$ (5)
+where $\Delta E_{ij}$ computes the International Commission on Illumination (CIE) color differences.
+B.3 Evaluation Metrics Beyond $\ell_{p}$ Norms
+LPIPS Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable] compares two images $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$ by measuring the distance between their deep feature representations extracted from a fixed CNN (e.g.
+AlexNet, VGG16, SqueezeNet) and linearly calibrating those distances with learned weights.
+$d(\boldsymbol{x}_{1},\boldsymbol{x}_{2})\;=\;\sum_{l}\frac{1}{H_{l}W_{l}}\sum_{h,w}\left\lVert\mathbf{w}_{l}\,\odot\!\bigl(\hat{y}^{\,l}_{1;h,w}-\hat{y}^{\,l}_{2;h,w}\bigr)\right\rVert_{2}^{2}$ (6)
+where, $d(\mathbf{x}_{1},\mathbf{x}_{2})$ represents the distance between the first image $\mathbf{x}_{1}$ and the second image $\mathbf{x}_{2}$ in the embedding space.
+The index $l$ denotes the layer in the network, while $H_{l}$ and $W_{l}$ are the height and width of the feature map at that layer.
+The spatial indices $h$ and $w$ iterate over the positions in the feature map.
+$\hat{y}^{l}_{1;h,w}$ and $\hat{y}^{l}_{2;h,w}$ denote the feature activations of the first and second images, respectively, at layer $l$ and spatial location $(h,w)$.
+$\mathbf{w}_{l}$ is a channel-wise weight vector or importance mask applied to the features, and $\odot$ indicates the element-wise (Hadamard) product.
+The expression $\|\cdot\|_{2}^{2}$ denotes the squared $\ell_{2}$ norm, measuring the weighted difference between the two feature representations.
+FID.
+Fréchet Inception Distance (FID) [heusel2017gans] measures the distributional differences between real and generated images in the feature space of a pretrained network (typically Inception-v3).
+Let $\mathcal{X}_{r}$ be the set of real images and $\mathcal{X}_{g}$ the set of generated images.
+The FID score is computed as:
+$\operatorname{FID}(\mathcal{X}_{r},\mathcal{X}_{g})=\left\lVert\boldsymbol{\mu}_{r}-\boldsymbol{\mu}_{g}\right\rVert_{2}^{2}+\operatorname{Tr}\left(\boldsymbol{\Sigma}_{r}+\boldsymbol{\Sigma}_{g}-2\left(\boldsymbol{\Sigma}_{r}\boldsymbol{\Sigma}_{g}\right)^{\tfrac{1}{2}}\right)$ (7)
+Here, $\boldsymbol{\mu}_{r}$ and $\boldsymbol{\Sigma}_{r}$ are the mean and covariance of the real images’ feature embeddings, while $\boldsymbol{\mu}_{g}$ and $\boldsymbol{\Sigma}_{g}$ are those of the generated images.
+The operator $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix.
+Table 7: Effectiveness of DASH with varying loss for the robust CIFAR-100 Cui2024 [cui2024decoupled] model.
+Loss ASR (↑) SSIM (↑) LPIPS (↓) FID (↓)
+Base JPEG TVM Ensemble NRP Avg.
+SSIM 100.00 99.90 100.00 99.32 99.61 99.77 94.43 0.0139 41.08
+LPIPS 100.00 100.00 100.00 99.41 99.80 99.84 92.34 0.0155 45.89
+B.4 Defense Strategies
+Below, we provide a brief overview of the defense mechanisms evaluated against our DASH framework, categorized into robust models and post-processing defenses.
+B.4.1 Robust Models
+Robust models are trained to be inherently resilient to adversarial examples.
+madry2018towards first proposed Adversarial Training, which injects adversarially perturbed inputs (with correct labels), typically generated via PGD attacks, into the training process to enhance robustness.
+TRADES [zhang2019theoretically] improves upon adversarial training by replacing the standard cross-entropy loss with a Kullback–Leibler (KL) divergence-based loss that better balances accuracy and robustness.
+Subsequent methods further refine these ideas, building on adversarial training and TRADES.
+For example, Decoupled Adversarial Learning [cui2024decoupled], Better Adversarial Training [wang2023better], Adversarial Weight Perturbation [bartoldson2024adversarial], and Efficient Robust Training [addepalli2022efficient] introduce modifications to improve training efficiency, robustness, or generalization.
+B.4.2 Post-Processing Defenses
+Post-processing defenses aim to increase the robustness of a given model (standard or robust) by transforming or purifying inputs at inference time.
+JPEG Compression [guo2017countering] applies a lossy compress–decompress operation before classification, which removes high-frequency components where adversarial noise typically resides.
+Total Variation Minimization (TVM) [wang2018adversarial] similarly aims to suppress high-frequency perturbations through optimization-based denoising.
+Neural Representation Purifier (NRP) [naseer2020self] is a more advanced approach that learns a purification network.
+Given a potentially adversarial image, this network iteratively adjusts pixels until a frozen reference model confidently recognizes the input as a clean example.
+NRP has demonstrated strong empirical performance and is widely adopted in recent adversarial defense research.
+Appendix C Base Attack Parameters and Their Selection Acorss Stages
+For the base attacks, we use Torchattacks [kim2020torchattacks], a PyTorch-based library that provides implementations of various adversarial attacks.
+The parameters for each attack used in our experiments are detailed below.
+For $\ell_{\infty}$-norm attacks, we set the perturbation budget to $\epsilon=\frac{7}{255}$.
+For different variants of FGSM evaluated in this paper, we use a momentum factor of 1.
+In TI-FGSM [dong2019evading] and DI-FGSM [xie2018improving], we set the resize_rate to 0.9, diversity_prob to 0.7, and random_start to False.
+In PI-FGSM++ [gao2020patch], we set the probability of using diverse inputs to 0.7 and the project_factor to 0.8.
+For the Fast Adaptive Boundary (FAB) attack [croce2020minimally], we set $\alpha_{\text{max}}=0.1$, overshooting factor $\eta=1.05$, and the backward step parameter $\beta=0.9$.
+For BIM [kurakin2016adversarial] and PGD [madry2017towards], we use a step size of $\alpha=0.01$ and number of iterations $T=10$.
+For the CW attack [carlini2017towards], we set the confidence parameter $c=10$, margin parameter $\kappa=0$, and optimize for 100 steps using a learning rate of $0.01$.
+C.1 Base Attack Selection Across Stages
+DASH learns the parameter values during training that are used to combine multiple base attacks at each stage of the adversarial example generation process.
+In Figure 5, we visualize the selected attack profiles for a robust model trained on CIFAR-100 [cui2024decoupled].
+The blue dotted line represents the initial combination of base attacks, while the orange line shows the learned combination after training.
+Initially, the CW attack is selected more heavily due to its strong misclassification and better perceptual preservation capability.
+However, as training progresses, DASH shifts weight away from CW, since it no longer contributes to the meta-loss—largely due to the zero-gradient problem introduced by the hinge loss (see Eq.
+(4)) with $\kappa=0$.
+As a result, our attack strategy continues to optimize for stronger and more confident adversarial examples, enhancing its transferability and effectiveness against various unseen post-processing defenses.
+This comes at the cost of slightly worse perceptual similarity scores, due to the increased distortion required to maintain high confidence.
+Appendix D DASH on Black-Box Attacks
+For measuring the performance of DASH with the Black-Box attack, we have taken 2 query-based attacks, Sign-Flip [chen2020boosting] and Rays [chen2020rays], and 2 decision-based attacks, Sign-Hunter [al2020sign] and Square attack [andriushchenko2020square], and combined them using DASH.
+They are the current state-of-the-art black-box attacks.
+Table 5 shows the performance comparison of DASH with other SOTA black-box attacks for 2 different models, ResNet-18 [he2016deep] and Vision Transformer [dosovitskiy2020image], trained on the CIFAR-10 dataset.
+The table shows that DASH achieves the highest average attack success rate for both of the models, exceeding the strongest baselines by almost 3% and 1%.
+It also maintains the highest imperceptibility with maximum SSIM and FID scores.
+This result implies that DASH is also useful for black-box attacks.
+Appendix E DASH with Varying Base Attacks
+Table 6 shows the DASH performance with varying base attacks.
+In all the combinations DASH shows an average attack success rate over 95% and SSIM over 90%, which implies that DASH is compatible with diverse base attacks and consistently achieves improved performance.
+In our main experiment, we used a meta-loss for optimizing attack weights in DASH, balancing attack success rate (ASR) and imperceptibility.
+Focusing on metrics aligned with the human visual system, we primarily report results using SSIM $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{ssim}}\cdot(1-\text{SSIM}(\boldsymbol{x},\boldsymbol{x}_{a}))$], but we also experimented with LPIPS $\mathcal{L}_{\text{total}}=\lambda_{\text{asr}}\cdot f_{y}(\boldsymbol{x}_{a})+\lambda_{\text{lpips}}\cdot\text{LPIPS}(\boldsymbol{x},\boldsymbol{x}_{a})$] and achieved similar output.
+Table 7 demonstrates that DASH  maintains a high ASR with minimal perturbation when optimizing with both SSIM and LPIPS.
+Specifically, the ASR varies by less than 0.1% and the SSIM score varies by only 0.02 units (assuming the 0 to 1 range), confirming the robustness of the framework regardless of the chosen perceptibility metric.